OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 5) with a slope of 2 in standard form.

OpenStudy (asadkarim7):

y=mx+c 5=2(-1) +c c= 7 therefore y=2x+7

OpenStudy (anonymous):

use y-y1 = m(x-x1) m=2 and x1,y1 = -1,5 so y-5 = 2(x-(-1)) y = 2x + 2 +5 2x - y = -7

OpenStudy (anonymous):

i'm from uk so i'm not sure what standsad form is in us ik uk its ax + by = c

OpenStudy (anonymous):

y = mx + b is also a "standard form", confusing I think..